The department of health studied the number of patients who need liver transplantation. The following data are the Liver Transplantation Waiting List (LTWL), where y is the size (in number of patients) and x is the corresponding year: y1375 1803 2772 386348416935 7635|7555 776684638412 8568 8921 x2001 2002 2003 2004 2005 2006 2007 2008 2009 2010 2011 2012 2013 Round your intermediate answers to four decimal places (e.g. 98.7654) (a) Fit the simple linear regression model using the method of least squares. Find the estimate of σ. Round your answer to the nearest integer (e.g. 9876) (b) How does the LTWL size change in average for a year? Round your answer to the nearest integer (e.g. 9876) βι (c) Estimate the number of patients in the list in 2017 (d) Calculate the fitted value of y corresponding to 2012. Find the corresponding residual. Round your answer to the nearest integer (e.g. 9876) Round your answer to the nearest integer (e.g. 9876)
Answer #1

a.

Standard Error of Y on X i.e Y = bo + b1 X

Xi Yi Y'=-1338520.84+669.95*X Y-Y' (Y-Yi)^2
2001 1375 2051.11 -676.11 457124.732
2002 1803 2721.061 -918.061 842836
2003 2772 3391.012 -619.012 383175.856
2004 3863 4060.963 -197.963 39189.349
2005 4841 4730.914 110.086 12118.927
2006 6935 5400.865 1534.135 2353570.198
2007 7635 6070.816 1564.184 2446671.586
2008 7555 6740.767 814.233 662975.378
2009 7766 7410.718 355.282 126225.3
2010 8463 8080.669 382.331 146176.994
2011 8412 8750.62 -338.62 114663.504
2012 8568 9420.571 -852.571 726877.31
2013 8921 10090.522 -1169.522 1367781.708

Standard error = Sqrt( ( sum ( Y -Yi )^2/ n-2 )

sum ( Y -Yi )^2 = 9679376.606

Standard Error = 9679376.606

variance = standard error^2 =93690331480780

b.

Line of Regression Y on X i.e Y = bo + b1 X
X Y (Xi - Mean)^2 (Yi - Mean)^2 (Xi-Mean)*(Yi-Mean)
2001 1375 36 22042303 28170
2002 1803 25 18206633 21335
2003 2772 16 10876297 13192
2004 3863 9 4870510 6621
2005 4841 4 1510252 2458
2006 6935 1 748358 -865
2007 7635 0 2449466 0
2008 7555 1 2205453 1485
2009 7766 4 2876677 3392
2010 8463 9 5726817 7179
2011 8412 16 5485324 9368
2012 8568 25 6240388 12490
2013 8921 36 8128639 17106

calculation procedure for regression

mean of X = sum ( X / n ) = 2007

mean of Y = sum ( Y / n ) = 6069.9231

sum ( (Xi - Mean)^2 ) = 182

sum ( (Yi - Mean)^2 ) = 91367117

sum ( (Xi-Mean)*(Yi-Mean) ) = 121931

b1 = sum ( (Xi-Mean)*(Yi-Mean) ) / sum ( (Xi - Mean)^2 )

= 121931 / 182

= 670

bo = sum ( Y / n ) - b1 * sum ( X / n )

bo = 6069.9231 - 670*2007 = -1338521

value of regression equation is, Y = bo + b1 X

Y'=-1338521+670* X

Answer:

beta1 =b1 =670

slope of equation =b1=670

c.

value of regression equation is, Y = bo + b1 X

Y'=-1338521+670* X

when X =2017 then

Y'=-1338521+(670* 2017)

Y' =12869

d.

Y'=-1338521+670* X

when X =2012 then

Y'=-1338521+(670*2012)

Y' =9519